OpenStudy (anonymous):

Evaluate x^2+2x if x=-2

OpenStudy (uri):

x^2+2x if x=-2 Put -2 where there is x So, (-2)^2+2(-2) Can you solve it?

OpenStudy (uri):

Wait it's -2^ no bracket.

OpenStudy (uri):

-2^2*

OpenStudy (anonymous):

not at all

OpenStudy (phi):

uri you were correct the first time: \[ x^2+2x \]if x=-2 Put -2 where there is x So, (-2)^2+2(-2) another way to think of it: factor out an x to get \[ x(x+2) \] sub in -2 for x: \[ -2(-2+2) \]

OpenStudy (uri):

Oh all right i was confused :) Thanks.

OpenStudy (phi):

@cynthia_franco uri told you the rule: everywhere you see x, erase the x and put in -2 you get \[ (-2)^2+2(-2) \] remember that (-2)^2 means -2 * -2, so this is the same as \[ -2 \cdot -2 + 2 \cdot -2 \] can you simplify that ?

OpenStudy (phi):

you do it in steps: first what is -2 times -2 ? then do 2 times -2 then add those two numbers